OpenStudy (anonymous):

The equation below shows the perimeter f(x) of several rectangular paintings having width x inches: f(x) = 40 + 2x The perimeter of each painting is less than 92 inches. Which of the following best describes the domain of f(x)?

OpenStudy (anonymous):

All real numbers greater than 0 and less than 26 All real numbers greater than 0 and less than 92 All integers greater than 0 and less than 26 All integers greater than 0 and less than 92

OpenStudy (alexandervonhumboldt2):

All real numbers greater than 0 and less than 92 i think

OpenStudy (alexandervonhumboldt2):

can yuo write a testimonial?

OpenStudy (anonymous):

could you help me with2 more @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (anonymous):

4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@DanJS